1. Suppose that a tax payer is in the 15% tax-rate bracket for the federal individual income tax and faces a 5% state income tax. a. If the taxpayer cannot deduct either tax against the other, what is the taxpayer’s combined marginal tax rate? What is the marginal tax rate if the taxpayer itemizes federal deductions and deducts the state tax? What if there was reciprocal deductibility? b. Now calculate all three combined marginal tax rates assuming that the state income tax rate is 10%. How do they change?
Answer #1

1.

Without deduction, combined tax rate = 15% + 5% = 20%

With state tax rate being deducted under federal tax, combined tax rate = 15% + 5% *(1-0.15) = 15% + 4.25 = 19.25%

With federal tax being deducted under state tax, combined tax rate = 5% + 15%*(1-0.05) = 5% + 14.25% = 19.25%

2.

Without deduction, combined tax rate = 15% + 10% = 25%

With state tax rate being deducted under federal tax, combined tax rate = 15% + 10% *(1-0.15) = 23.5%

With federal tax being deducted under state tax, combined tax rate = 10% + 15%*(1-0.1) = 23.5%

Combined tax rate increased in second case